Some batteries have lithium and cobalt in them. The Li0 gets oxidized to Li1+, and Co2+ gets reduced to Co0.

Write a balanced reaction for Li oxidation and Co reduction. Make sure the stoichiometry and charge are balanced!
Answer #1

Reduction of Co2+(aq)

Co2+(aq) + 2e = Co(s)

Oxidation of lithium

Li(s) = Li+(aq) + e ( multiply by 2 )

2 Li(s) = 2 Li(s) + 2e

Add them , we get

2 Li(s) + Co2+(aq) = Co(s) + 2 Li+(aq)

If you have any questions please comment

If you satisfied with the solution please rate it thanks